OpenStudy (anonymous):

How would you graph this? My original equation= F(x) = 23(1.2)^x A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). It would be like 1+ -6, 2+ -6 and so on you just substitute it into the equation to make any profit he would have to do more than 6. F(x) - 6= 23( 1.2^x) - 6

OpenStudy (jdoe0001):

so... what area you trying to do? f(x) + g(x) or graph \(\bf f(x) = 23(1.2)^x\) ?

OpenStudy (anonymous):

it said to graph F(x) + G(x) and im not sure how to do that

OpenStudy (jdoe0001):

so... ok.. what's f(x) and what's g(x) firstly ?

OpenStudy (anonymous):

it is -6

OpenStudy (jdoe0001):

what do you mean? hmm if you dunno what f(x) and g(x) are ..... kinda hard to add them up :) I can see the g(x) = -6 up there, what about f(x) ? is it \(\bf f(x) = 23(1.2)^x?\)

OpenStudy (anonymous):

I think you just substitute a number but i'm not sure, if x was one than it would be F(-5) because your subtracting 6 right?

OpenStudy (jdoe0001):

if g(x) = -6 and \(f(x) = 23(1.2)^x\) then \(\bf f(x) - 6= 23( 1.2^x) - 6\) is correct

OpenStudy (anonymous):

but im not sure how you would simplify it to be able to graph it,

OpenStudy (jdoe0001):

well... you can;t simplify it really

OpenStudy (anonymous):

Oh wait so it says just graph f(x) + g(x) so how would you graph F(x) - 6

OpenStudy (jdoe0001):

what you can do is keep in mind that the graph, is just the graph of \(\bf 1.2^x\) only and is "shrunk" by a factor of 23, and shifted downwards by 6 units

OpenStudy (anonymous):

so it would end up being on 17 with the same rate of change?

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards \\ \quad \\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ % template start y = {\color{purple}{ 23}} ( 1.2^{{\color{blue}{ B}}x + {\color{red}{ C}}} ) - {\color{green}{ D}}\\ % template ends \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to\ the\ left \\ \quad \\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to\ the\ right\end{array} \end{array} }\)

OpenStudy (anonymous):

i think its supposed to be a verticle shift and just move down 6 to make the principle 16?

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

so yes

OpenStudy (anonymous):

Thanks

OpenStudy (jdoe0001):

yw